OpenStudy (anonymous):

solving systems of equations by elimination -x + 9y = -5 x - 5y = 1

OpenStudy (haleyelizabeth2017):

Add the second equation to the first equation to get 4y=-4

OpenStudy (haleyelizabeth2017):

Then, divide -4 by 4 to get y=-1. Plug in -1 for y into either equation and solve for x.

OpenStudy (haleyelizabeth2017):

x-5y=1 x-5(-1)=1 x+5=1 x=-4

OpenStudy (haleyelizabeth2017):

Then plug x and y back into either equation to check your answer.

OpenStudy (anonymous):

do i add or subtract them

OpenStudy (haleyelizabeth2017):

add :)

OpenStudy (anonymous):

ty :D

OpenStudy (haleyelizabeth2017):

No problem!

OpenStudy (anonymous):

-4+-1

OpenStudy (anonymous):

with will be -5

OpenStudy (anonymous):

whitch*

OpenStudy (haleyelizabeth2017):

Wait..

OpenStudy (haleyelizabeth2017):

-(-4) + 9(-1) = -5 (-4) - 5(-1) = 1 Plug them into these equations to check

OpenStudy (anonymous):

omg im new too this subject lol like ugh

OpenStudy (haleyelizabeth2017):

haha don't worry about it, you'll get the hang of it!

OpenStudy (haleyelizabeth2017):

just know...the solution would be (-4, -1)

OpenStudy (anonymous):

for the answer right ill let you kno if i need help on the next one

OpenStudy (haleyelizabeth2017):

Okay

OpenStudy (haleyelizabeth2017):

Did you understand how I got that?

OpenStudy (anonymous):

thanks again

OpenStudy (haleyelizabeth2017):

No problem!

OpenStudy (anonymous):

y = 2/3x-1 y=3

OpenStudy (anonymous):

y = 2/3x-1 y=3

OpenStudy (haleyelizabeth2017):

are those the two equations given?

OpenStudy (anonymous):

yes

OpenStudy (haleyelizabeth2017):

okay, so 3=2/3x-1 y is given so we have (_, 3) so far.

OpenStudy (anonymous):

ok

OpenStudy (haleyelizabeth2017):

3+1=2/3x 4=2/3x 4*3/2=x x=6 so (6, 3) :)

OpenStudy (anonymous):

THANKS!

OpenStudy (haleyelizabeth2017):

No problem!

OpenStudy (anonymous):

8 more to go idk if you going to help me with all this lmao